# Wormhole-Induced correlation: A Link Between Two Universes

### v2

@@ Abstract
Motivated by the profound connection between quantum mechanics and spacetime geometry, particularly the conjectured correspondence between wormholes and quantum entanglement as proposed in the ER=EPR framework, this study aims to investigate the influence of wormhole geometries on quantum information extraction. We examine the correlation-specifically mutual information (MI) and entanglement-extracted by two Unruh-DeWitt (UDW) detectors from the quantum vacuum field in the presence of a BTZ wormhole featuring a null-like throat, also known as an Einstein-Rosen bridge. First, we analyze how the detector’s position relative to the wormhole throat and the throat’s size affect the extracted MI. Our results indicate that the wormhole enhances MI extraction, with maximal MI achieved when the detectors are located at specific image-symmetric points connected by the wormhole. By analyzing the behavior of the nonlocal contribution term and the classical noise term, it is found that the correlations extracted contain genuine non-classical components. This work highlights the feasibility of extracting quantum correlations through null-like wormhole geometries and provides a novel perspective for probing the potential relationship between spacetime topology and the nonlocal characteristics of quantum mechanics.
@@
To explore the interplay between wormhole geometry and quantum nonlocality, this study investigates the harvesting of quantum correlations in a spacetime containing a BTZ wormhole with a null-like throat. Specifically, we employ a pair of UDW particle detectors to analyze how the throat size of the wormhole and the spatial arrangement of the detectors affect the extracted correlations. Our results show that the wormhole structure acts as a geometric bridge enabling the emergence of quantum correlations between classically disconnected regions. The harvested correlations exhibit both local and nonlocal components, with the both enhanced as the detectors approach symmetric positions relative to the throat. Notably, when the throat size approaches zero—effectively in the absence of a wormhole—these correlations vanish, underscoring their geometric origin. This work demonstrates the feasibility of extracting quantum correlations through wormhole geometries and provides a concrete framework for probing the relationship between spacetime topology and quantum nonlocality.
@@ II.1 BTZ Wormhole
$U=\pm\frac{e^{-\kappa u}}{\kappa},\quad V=\pm\frac{e^{\kappa v}}{\kappa},$ (7)
@@
Figure 2: The Penrose diagram of the extended BTZ spacetime. Regions I and IV as two asymptotically flat universes connected via an event horizon. Regions II and III represent the interiors of the black hole and white hole, respectively. The wavy lines at the top and bottom denote the singularity, while special spacelike hypersurfaces (U=VU=-V or $T=0$) traversing both universes are indicated by yellow dashed lines.
@@
$W_{wh}(\mathrm{x,x^{\prime}})=\frac{1}{4\pi\sqrt{2}l}\sum^{+\infty}_{n=-\infty}\left[\frac{1}{\sqrt{\sigma_{\epsilon}(\mathrm{x,\Gamma^{n}x^{\prime}})}}-\frac{\xi}{\sqrt{\sigma_{\epsilon}(\mathrm{x,\Gamma^{n}x^{\prime}})}}\right],$ (11)
@@ III Result
The wormhole size does not significantly impact the MI peak value but does affect the shape of the peak region. Notably, in this asymmetric distribution, we find that the MI vanishes when detector Alice is located at the wormhole throat. This can be attributed to the extreme relative Hawking temperature [30]. Overall, this peak region suggests that the wormhole bridges two special image-symmetrical locations, characterized by an optimal correlation.
@@
In summary, we employ the UDW detectors model to investigate how the size of the wormhole throat and the distribution of detectors influence the extracted correlation. Our results indicate that the presence of a wormhole, characterized by a nonzero throat size, enables detectors located on opposite sides to extract MI. Furthermore, when detector Bob is fixed, we observe a peak in the extracted MI, with the maximum occurring when Alice is positioned symmetrically. Notably, the extracted MI remains nonzero as long as the throat size is nonzero; although it diminishes rapidly as the distance between the detectors and the throat increases. These peaks suggest that the wormhole acts as a conduit between two specific symmetrical locations, resulting in the MI extracted between them reaching an optimal value. Subsequently, by analyzing the behavior of the nonlocal contribution term and the classical noise term, as well as by computing the concurrence between the two detectors, we demonstrate that the harvested correlations contain nonlocal correlations. These findings demonstrate that the existence of a null-like wormhole can enhance quantum correlation behaviors between classically causally disconnected regions.

### v3

@@ Abstract
Motivated by the profound connection between quantum mechanics and spacetime geometry, particularly the conjectured correspondence between wormholes and quantum entanglement as proposed in the ER=EPR framework, this study investigate the influence of wormhole geometries on quantum information extraction. We examine the correlation-specifically mutual information (MI) and entanglement-extracted by two Unruh-DeWitt (UDW) detectors from the quantum vacuum field in the presence of a Bañados-Teitelboim-Zanelli (BTZ) wormhole featuring a null-like throat, also known as an Einstein-Rosen bridge. First, we analyze how the detector’s position relative to the wormhole throat and the throat’s size affect the extracted MI. Our results indicate that the wormhole enhances MI extraction, with maximal MI achieved when the detectors are located at specific image-symmetric points connected by the wormhole. By analyzing the behavior of the nonlocal contribution term and the classical noise term, it is found that the correlations extracted contain genuine non-classical components. This work highlights the feasibility of extracting quantum correlations through null-like wormhole geometries and provides a novel perspective for probing the potential relationship between spacetime topology and the nonlocal characteristics of quantum mechanics.
@@
To explore the interplay between wormhole geometry and quantum nonlocality, this study investigates the harvesting of quantum correlations in a spacetime containing a Bañados-Teitelboim-Zanelli (BTZ) wormhole with a null-like throat. Specifically, we employ a pair of UDW particle detectors to analyze how the throat size of the wormhole and the spatial arrangement of the detectors affect the extracted correlations. Our results show that the wormhole structure acts as a geometric bridge enabling the emergence of quantum correlations between classically disconnected regions. The harvested correlations exhibit both local and nonlocal components, with the both enhanced as the detectors approach symmetric positions relative to the throat. Notably, when the throat size approaches zero—effectively in the absence of a wormhole—these correlations vanish, underscoring their geometric origin. It should be underscored that these correlation harvesting phenomena do not imply any possibility of communication between the two universes, as classical signals cannot traverse the wormhole.
@@ II.1 BTZ Wormhole
$U=\pm\frac{e^{-\kappa\tilde{u}}}{\kappa},\quad V=\pm\frac{e^{\kappa\tilde{v}}}{\kappa},$ (7)
@@
Figure 2: The Penrose diagram of the extended BTZ spacetime. Regions I and IV as two universes connected via an event horizon. Regions II and III represent the interiors of the black hole and white hole, respectively. The wavy lines at the top and bottom denote the spacelike infinity ($r=0$), while special spacelike hypersurfaces (U=VU=-V or $T=0$) traversing both universes are indicated by yellow dashed lines.
@@
$W_{wh}(\mathrm{x,x^{\prime}})=\frac{1}{4\pi\sqrt{2}l}\sum^{+\infty}_{n=-\infty}\left[\frac{1}{\sqrt{\sigma_{\epsilon}(\mathrm{x,\Gamma^{n}x^{\prime}})}}-\frac{\xi}{\sqrt{\sigma_{\epsilon}(\mathrm{x,\Gamma^{n}x^{\prime}})+2}}\right],$ (11)
@@ III Result
The wormhole size does not significantly impact the MI peak value but does affect the shape of the peak region. Notably, within this asymmetric configuration, the MI progressively diminishes as detector Alice approaches the wormhole throat. This can be attributed to the extreme relative Hawking temperature [30]. Overall, this peak region suggests that the wormhole bridges two special image-symmetrical locations, characterized by an optimal correlation.
@@
In summary, we employ the UDW detectors model to investigate how the size of the wormhole throat and the distribution of detectors influence the extracted correlation. Our results indicate that the presence of a wormhole, characterized by a nonzero throat size, enables detectors located on opposite sides to extract MI. Furthermore, when detector Bob is fixed, we observe a peak in the extracted MI, with the maximum occurring when Alice is positioned symmetrically. Notably, the extracted MI remains nonzero as long as the throat size is nonzero; although it diminishes rapidly as the distance between the detectors and the throat increases. These peaks suggest that the wormhole acts as a conduit between two specific symmetrical locations, resulting in the MI extracted between them reaching an optimal value. Subsequently, by analyzing the behavior of the nonlocal contribution term and the classical noise term, as well as by computing the concurrence between the two detectors, we demonstrate that the harvested correlations contain nonlocal correlations. These findings demonstrate that the existence of a null-like wormhole can enhance quantum correlation behaviors between classically causally disconnected regions. It is essential to emphasize, nevertheless, that such correlation harvesting phenomena do not enable inter-universe communication, since classical signals remain unable to traverse the wormhole.